On Tuesday the Beef Market sold 400 pounds of prime rib steak at $9.98 per pound and 120 pounds of rib-eye steak at $6.49 per pound. What was the average cost in dollars per pound of the steaks sold on Tuesday? Express your answer to the nearest hundredth.
Answer by Theo(13342)

400 * 9.98 + 120 * 6.49 = 4770.8 total cost for steaks (both rib and rib-eye).
400 + 120 = 520 pounds total.
4770.8 / 520 = 9.174615385 per pound for steaks (both rib and rib-eye)
that's the overall average cost per steak.
that's 9.17 per pound expressed to the nearest hundredth.